1.Consider a hypothetical ionic compound AB (comprised of At and B ions). Given the following enthalpy data and using a Born-Ha2. Rank the following ionic compounds in order of lattice energy, from the least exothermic (least negative lattice enthalpy) to
Answer #1

SHSub & TE+ EA + V -422= 105+ 109 + 530 - 379+ U U= [-787) Klimos ③ Rbl. Cao, sro, NaCl Rbl < Nacik sro <GO